OpenStudy (anonymous):

is 25x^2+40x+64 a perfect square trinomial?

OpenStudy (anteater):

A perfect square trinomial follows this pattern: a^2 + 2ab + b^2 First ask yourself, what are a and b for your trinomial?

OpenStudy (anteater):

That is, what would you "square" to get the first term, 25x^2? And what would you "square" to get the last term, 64?

OpenStudy (anteater):

So then, 25x^2 = (5x)^2 so a = 5x 64 = 8^2 so b = 8 In order for this to be a perfect square trinomial, 40x must equal 2ab; 2ab = 2(5x)(8) = 80x ... so this is not a perfect square trinomial. But I have a feeling you already determined that :D